OpenStudy (anonymous):

whos in algebra 2????

OpenStudy (anonymous):

@puledo_cat @Ghostboy2872 @Aya713

OpenStudy (anonymous):

algebra 1, sorryyyyyyy only 9th grade class i have

OpenStudy (anonymous):

its ok. im a ninth grader but ive already completed algebra 1 and geometry so im stuck in algebra 2

OpenStudy (anonymous):

dang lucky^

OpenStudy (anonymous):

i can help you out though my sister is in algebra 1 as well and i help her. what school do you go to?

OpenStudy (anonymous):

I have integrated math 2

OpenStudy (anonymous):

do you think you can help me?? @puledo_cat

OpenStudy (anonymous):

depends what the question is but I'll give it a try

OpenStudy (anonymous):

what is an equation of the line with slope 3 that goes through point (-1/2,5/2)?

OpenStudy (anonymous):

so an equation for this is y=mx+b but we have to find out the constant which is b and we already have the slope which is m.

OpenStudy (anonymous):

the slope is 3 so y=3x+b, but now we need to find out what b equals

OpenStudy (anonymous):

wouldnt b equal to 4?

OpenStudy (anonymous):

why do you think it equals 4?

OpenStudy (anonymous):

its just a saying i got taught if m is equal to 3 then b is equal to 4

OpenStudy (anonymous):

i think the answer is d

OpenStudy (anonymous):

oh what you need to do is y=-1/2 and x=5/2 so -1/2=3(5/2)+b

OpenStudy (anonymous):

now we have to solve for b

OpenStudy (anonymous):

so -1/2=7.5??

OpenStudy (anonymous):

nope, first you have the equation right which is -1/2=3(5/2)+b so we start off by multiplying 3*5/2

OpenStudy (anonymous):

which equals 15/2 because 3 times 5 is 15 and 2 times 1 is 2 so 15/2

OpenStudy (anonymous):

and 15/2 is 7.5

OpenStudy (anonymous):

is equal****

OpenStudy (anonymous):

yes but in this case we can use it as a fraction. After that we subtract -15/2 to both sides and you get -16/2

OpenStudy (anonymous):

so what is -16 divided by 2?

OpenStudy (anonymous):

-8?

OpenStudy (anonymous):

yes so b is equaled to -8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but thats not an answer choice

OpenStudy (anonymous):

so y=3x-8

OpenStudy (anonymous):

thats not an answer choice

OpenStudy (anonymous):

its ok i got the answer

OpenStudy (anonymous):

alright sorry I tried my best ;p